UT 20160422 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson,
Observer AO: DMiller, GTaylor (both LBTO), EPinna, APuglisi, SEsposito (remote from Italy)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Startup AO-system, check NCPA, opt. gain etc on sky

- Depending on how and how long it goes, we may scwitch to LUCI1 AO-com than

Summary

- All night tests done by AO-group

- System seems to be in good shape. Optical gain needs some (minor) improvement

- Bin3 + 4 measurements are missing but are short and may be carried out tomorrow

Details

All times are UT
Files are luci1.20160422.NNNN.fits

01:59 Sunset, Steve opens, quite some Cirrus around, T = +9deg, Hum = 17%, low wind

02:20 Steve checks pointing and collimates, seeing around 1"

02:40 Doug takes over for AO-tests

Log at http://wiki.lbto.org/bin/view/AdaptiveOptics/20160421_AO_SX

04:30 Tests ongoing, DIMM-seeing 1.1-1.5", still lots of Cirrus

06:30 Still testing, DIMM-seeing 1.1-1.5", still lots of Cirrus

08:30 Still testing, DIMM-seeing 1.1-1.5", still lots of Cirrus, 1-2mag of extinction, wind about 10m/s

10:30 Still testing, DIMM-seeing 0.8-1.2", Cirrus less dense, wind about 7m/s

12:15 Its getting too bright, Steve's closing up

-- %USERSIG{JochenHeidt - 2016-04-21}%
Topic revision: r6 - 22 Apr 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback